A population of size N=150 has u=8 and standard deviation of o=5.4. What is the probability that a random sample of size n=20 will have a mean of 9.5 above.

1
Expert's answer
2022-05-17T00:16:36-0400

Given,

Mean of population "\mu=8"

Standard deviation "\sigma = 5.4"

Using

"P(\bar X >9.5)= \dfrac{\bar X-\mu}{\frac{\sigma}{\sqrt n}}"


"P( \bar X>9.5)= \dfrac{9.5-8}{\frac{5.4}{\sqrt{20}}}=\dfrac{1.5}{1.20}=1.25"


So, "P(Z>1.25)=0.1056"